OpenStudy (anonymous):

Find the distance between the two points. Round to the nearest tenth if necessary. (–4, –12), (11, –6)

OpenStudy (anonymous_user):

You'll get better help is you post your question in the Mathematics section. http://openstudy.com/groups/mathematics

OpenStudy (texaschic101):

distance formula : d = sq rt ((x2 - x1)^2 + (y2 - y1)^2)) (-4,-12).....x1 = -4 and y1 = -12 (11, -6)...x2 = 11 and y2 = -6 now we sub d = sq rt ((11 - (-4))^2 + (-6 - (-12))^2 d = sq rt ((11 + 4)^2 + (-6 + 12)^2 d = sq rt (15^2) + (6^2) d = sq rt (225 + 36) d = sq rt 261 d = 16.155 rounded = 16.2 http://www.purplemath.com/modules/distform.htm

OpenStudy (anonymous):

-80